OpenStudy (anonymous):

write the slope intercept equation for the line that passes thru. (-4,-3) and (-6,7)

OpenStudy (dumbcow):

y = -5x-23

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help please and thanks

OpenStudy (anonymous):

m= (-3-7)/(-4-(-6) -10/2 = -5 y-7= -5(x-(-6) = -11/x -30+ 7 y = -5x-23

OpenStudy (anonymous):

sorry I tye wrong m= (-3-7)/(-4-(-6) -10/2 = -5 y-7= -5(x-(-6) = -5/x -30+ 7 y = -5x-23